The magnetic flux through each of five faces of a die (singular of ''dice'') is given by ΦB = ±N Wb, where N (= 1 to 5) is the number of spots on the face. The flux is positive (outward) for N even and negative (inward) for N odd. What is the flux through the sixth face of the die?
Answer #1

Let flux through sixth face is Wb

Total magnetic flux = 0

3 Wb ...... Answer
Answer #2

Calculating the Flux Through the Sixth Face of the Die

Given:

  • A die has 6 faces, labeled with N=1 to 6 spots.

  • Magnetic flux through five faces is given by:

    • ΦB=+NWb (outward) if N is even.

    • ΦB=NWb (inward) if N is odd.

  • The flux through the sixth face (unknown) is denoted as Φ6.

Key Principle: Gauss's Law for Magnetism

BdA=0

This means the total magnetic flux through a closed surface (like the die) must be zero.



Step 1: Calculate Flux Through the Five Known Faces

Face (N)Flux (ΦB)Reason
11WbN=1 is odd.
2+2WbN=2 is even.
33WbN=3 is odd.
4+4WbN=4 is even.
55WbN=5 is odd.
Total (5 faces)3WbSum: 1+23+45=3.


Step 2: Solve for the Sixth Face (N=6)

Using Gauss's Law:

Φtotal=Φ5 faces+Φ6=03Wb+Φ6=0    Φ6=+3Wb

Verification:

  • The sixth face has N=6 (even), so by the given rule:

    Φ6=+6Wb(if it followed the pattern)

  • However, Gauss's Law overrides this pattern, requiring the total flux to be zero. Thus:

    Φ6=+3Wb


Answer:

The flux through the sixth face is +3Wb (outward).


answered by: anonymous